OpenStudy (anonymous):

Dana plans to cover this model of a half moon completely with silver foil. ( imagine a semicircle with a radius of 8 cm) How many square centimeters of silver foil would Dana need to cover the model completely?

OpenStudy (anonymous):

I got 402 but it is wrong

OpenStudy (anonymous):

100,48

OpenStudy (anonymous):

circle area:\[A=\pi*r ^{2}\] since it is half moon: devide by 2

OpenStudy (anonymous):

is that 100.48?

OpenStudy (anonymous):

taking value for pi 3,14, yes

OpenStudy (anonymous):

but it said... cover the entire thing? the outside not the volume

OpenStudy (anonymous):

i understand it like if you would cut the chape of half moon from sheet of paper, and cover one of it's sides

OpenStudy (anonymous):

gotcha. can you look at this next one as well

jimthompson5910 (jim_thompson5910):

SA = (half of outside) + (inside circular surface) SA = (1/2)*pi*r^2 + pi*r^2 SA = (1/2)*4*pi*8^2 + pi*8^2 SA = 128*pi + 64*pi SA = 192*pi So the exact surface area is 192pi square cm and it approximates to 603.18578948924 square cm (use a calculator)